Help me please! A coffee and tea shop wishes to blend a $3.50-per-pound coffee with a $4.75-per-pound coffee in order to produce a $4-per-pound blend. How much of each would hve to be used to produce 100 pounds of the nw blend?
Answer by ptaylor(2198)

Let x=amount of 3.50 coffee
Then 100-x=amount of 4.75 coffee
Now we know that the total cost of the 3.50 coffee (3.50x) plus the total cost of the 4.75 coffee(4.75)(100-x) must equal the total cost of the final mixture (4.00(100)). So our equation to solve is:
3.50x+4.75(100-x)=4.00(100) get rid of parens
3.50x+475-4.75x=400 subtract 475 from both sides
3.50x+474-475-4.75x=400-475 collect like terms
-1.25x=-75 divide both sides by -1.25
x=60 lbs -------------------------------amount of 3.50 coffee
100-x=100-60=40 lb------------------------amount of 4.75 coffee
CK
3.50(60)+4.75(40)=4.00(100)
210+190=400
400=400

Hope this helps----ptaylor